OpenStudy (anonymous):

36/99 divided by 44/99

jimthompson5910 (jim_thompson5910):

`36/99 divided by 44/99` is the same as `36/99 times 99/44` notice how the second fraction flipped and I changed the division to multiplication

OpenStudy (anonymous):

so you turn the second fraction to the recipricol and change the equation to multiplication

OpenStudy (anonymous):

whenever you divide or is there another way?

jimthompson5910 (jim_thompson5910):

yes because of this rule \[\Large \frac{A}{B} \div \frac{C}{D} = \frac{A}{B} \times \frac{D}{C}\]

OpenStudy (anonymous):

what is that rule called?

jimthompson5910 (jim_thompson5910):

I'm not sure it has a specific name. Let me look

OpenStudy (anonymous):

so Idk how to multiply this new equation..?

jimthompson5910 (jim_thompson5910):

ok so you understand how we have gone from \[\Large \frac{36}{99} \div \frac{44}{99}\] to \[\Large \frac{36}{99} \times \frac{99}{44}\] right? or no?

jimthompson5910 (jim_thompson5910):

here's another example http://www.coolmath.com/sites/cmat/files/images/1_3dividedby4_5.gif

jimthompson5910 (jim_thompson5910):

Notice we have this pair of 99s \[\Large \frac{36}{{\color{red}{99}}}\times\frac{{\color{red}{99}}}{44}\] they will cancel (since one 99 is up top, the other 99 is down below) \[\Large \frac{36}{{\color{red}{99}}}\times\frac{{\color{red}{99}}}{44} = \frac{36}{{\color{red}{\cancel{99}}}}\times\frac{{\color{red}{\cancel{99}}}}{44}\] \[\Large \frac{36}{{\color{black}{99}}}\times\frac{{\color{black}{99}}}{44} = \frac{36}{1}\times\frac{1}{44}\] \[\Large \frac{36}{{\color{black}{99}}}\times\frac{{\color{black}{99}}}{44} = \frac{36 \times 1}{1 \times 44}\] \[\Large \frac{36}{{\color{black}{99}}}\times\frac{{\color{black}{99}}}{44} = \frac{36}{44}\] making sense so far?

OpenStudy (anonymous):

yes thank you